October 3, 2022

From HallCWiki
Jump to: navigation, search

General Update:

  • Beam recovery to Hall C around 10 am on Saturday.
  • Problem with the high current beam delivery due to the dump viewer ion chamber trips at about 30 uA on LH2 target.
    • New IC calibration on LH2 didn't help.
    • Harp scans and hole check looked fine.
    • IC FSD card spare that was used for the replacement during last week's maintenance day was not a good one. It was swapped one more time and the problem was fixed --> good 65 uA beam.
  • Finished the coincidence elastic data taking including some sieve runs. Also, completed the coincidence elastic with the SHMS delta scan.
  • During early Sunday owl shift, switched from COIN DAQ to SINGLES DAQ (Casey Morean + Bob Michaels). Took SHMS/HMS Q2 defocused data.

Problems resolved:

  • Sunday swing: SHMS HB interlock with fast discharge (engineer on call helped reseting it)
  • Monday owl: HMS rotation emergency stop (engineer on call came in to reset the drive controller)
  • Missing plots on the online analysis when switched to single arm running. https://logbooks.jlab.org/entry/4050289

Things need to be fixed:

  • Ramping up the SHMS to 9.005 GeV yesterday, we found out that the SHMS Q2 would stop before reaching the MOL (SHMS Q2 setting is 3475.9 but it stops ramping at ~3415A).
  • SHMS NGCER HV (Dave Gaskell and Mark Jones modified the HV on PMT 1, will check it with the real data once we get beam back!)